# Mean-field games with rough common noise I: the linear-quadratic case

### v1

@@ 3.1 Assumptions, state dynamics and cost functional
From this point on (and throughout the rest of the paper), we assume that $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in\mathbb{R}_{+}},\mathbb{P})$ is the usual augmentation of a filtration generated by a $q$-dimensional Brownian motion $W$ and a $\sigma$-algebra $\mathcal{F}_{0}^{0}$ independent of WW444this assumption is made to avoid technicalities with martingale representation arguments that are used in the derivation of the Volterra FBSDEs. From now on and for the rest of this section, we will work under the following assumptions:
@@
over all admissible controls $\alpha\in\mathcal{A},$ where
@@ 3.2 The stochastic maximum principle
$=R^{m}_{1}(X^{\alpha,m,\xi},(\delta X)^{h},\alpha,\varepsilon h)+R^{m}_{2}((\delta X)^{h},\varepsilon h)$
$=\varepsilon R^{m}_{1}(X^{\alpha,m,\xi},(\delta X)^{h},\alpha,h)+\varepsilon^{2}R^{m}_{2}((\delta X)^{h},h).$
@@
for some constant $C^{\prime\prime}$ depending on $\mathcal{K},A,B,T,Q_{t},\bar{Q}_{t},R_{t},Q,\bar{Q}.$ Hence
@@
$\widehat{Y}^{\alpha,m,\xi}_{t}=\mathcal{K}(T,0)^{\top}\big((Q+\bar{Q})X_{T}^{\alpha,m,\xi}-\bar{Q}Sm_{T}\big)$
$\quad+\int_{t}^{T}\mathcal{K}(s,0)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{\alpha,m,\xi}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}\mathcal{K}(0,s)^{\top}\widehat{Y}^{\alpha,m,\xi}_{s}\Big)\mathrm{d}s-\int_{t}^{T}\widehat{Z}^{\alpha,m,\xi}_{s}\mathrm{d}W_{s}.$
The latter is a classical affine backward SDE with a unique solution $(\widehat{Y}^{\alpha,m,\xi},\widehat{Z}^{\alpha,m,\xi})\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{H}^{2}([0,T];\mathbb{R}^{d\times q})$, see for example [19, Theorem 4.23], hence, for every $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0}),\alpha\in\widehat{\mathcal{A}}$ and $m\in C([0,T];\mathbb{R}^{d})$, (3.2) also has a unique solution $({Y}^{\alpha,m,\xi},{Z}^{\alpha,m,\xi})\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{H}^{2}([0,T];\mathbb{R}^{d\times q})$.
@@
$(Y^{\alpha,m,\xi}_{T})^{\top}$ ${(\delta X)}_{T}^{h}=\big(\mathcal{K}(T,0)^{\top}Y^{\alpha,m,\xi}_{T}\big)^{\top}\big(\mathcal{K}(0,T){(\delta X)}_{T}^{h}\big)$
$=\int_{0}^{T}\big(\mathcal{K}(t,0)^{\top}Y^{\alpha,m,\xi}_{t}\big)^{\top}\mathrm{d}\big(\mathcal{K}(0,t){(\delta X)}_{t}^{h}\big)+\int_{0}^{T}\big(\mathcal{K}(0,t){(\delta X)}_{t}^{h}\big)^{\top}\mathrm{d}\big(\mathcal{K}(t,0)^{\top}Y^{\alpha,m,\xi}_{t}\big).$
@@
$\int_{0}^{T}\big(\mathcal{K}(0,t){(\delta X)}_{t}^{h}$ $\big)^{\top}\mathrm{d}\big(\mathcal{K}(t,0)^{\top}Y^{\alpha,m,\xi}_{t}\big)=-\int_{0}^{T}\Big(\big({(\delta X)}_{t}^{h}\big)^{\top}A_{t}^{\top}Y^{\alpha,m,\xi}_{t}\Big)\mathrm{d}t$
@@
$\mathbb{E}\Big[(Y^{\alpha,m,\xi}_{T})^{\top}{(\delta X)}_{T}^{h}\Big]=-\mathbb{E}\!\left[\int_{0}^{T}\big({(\delta X)}_{t}^{h}\big)^{\top}\big((Q_{t}+\bar{Q}_{t})X_{t}^{\alpha,m,\xi}-\bar{Q}_{t}S_{t}m_{t}\big)\mathrm{d}t\right]$
$+\mathbb{E}\!\left[\int_{0}^{T}\Big((Y^{\alpha,m,\xi}_{t})^{\top}B_{t}h_{t}\Big)\mathrm{d}t\right].$
@@
$\left\{\begin{aligned} &{X}^{m,\xi}_{t}=\mathcal{K}(t,0)\xi+M(m)_{t}\\ &\quad\quad+\int_{0}^{t}\mathcal{K}(t,s)\big(A_{s}{X}^{m,\xi}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}Y^{m,\xi}_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\mathcal{K}(t,s)\Sigma_{s}\mathrm{d}W_{s},\\ &Y^{m,\xi}_{t}=\mathcal{K}(T,t)^{\top}\big((Q+\bar{Q})X_{T}^{m,\xi}-\bar{Q}Sm_{T}\big)\\ &\quad\quad+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{m,\xi}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}Y^{m,\xi}_{s}\Big)\mathrm{d}s-\int_{t}^{T}\mathcal{K}(s,t)^{\top}Z^{m,\xi}_{s}\mathrm{d}W_{s}.\end{aligned}\right$. (17)
@@
$\left\{\begin{aligned} &{X}^{\alpha,m,\xi}_{t}=\mathcal{K}(t,0)\xi+M(m)_{t}\\ &\quad\quad+\int_{0}^{t}\mathcal{K}(t,s)\big(A_{s}{X}^{\alpha,m,\xi}_{s}+B_{s}\alpha_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\mathcal{K}(t,s)\Sigma_{s}\mathrm{d}W_{s},\\ &Y^{\alpha,m,\xi}_{t}=\mathcal{K}(T,t)^{\top}\big((Q+\bar{Q})X_{T}^{\alpha,m,\xi}-\bar{Q}Sm_{T}\big)\\ &\quad\quad+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{\alpha,m,\xi}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}Y^{\alpha,m,\xi}_{s}\Big)\mathrm{d}s-\int_{t}^{T}\mathcal{K}(s,t)^{\top}Z^{\alpha,m,\xi}_{s}\mathrm{d}W_{s}.\end{aligned}\right$.
@@
System (17), derived by (16) as explained in 3.5, a priori describes a necessary condition for $J^{m,\xi}(\cdot)$ to have a local minimizer in $\widehat{\mathcal{A}}$. Below we show that in fact it provides a necessary and sufficient condition for $J^{m,\xi}(\cdot)$ to have a unique global minimizer in $\widehat{\mathcal{A}}$, as in the non-Volterra setting.
@@
Thus, modulo well-posedness of the Volterra FBSDE (17), we have showed following:
@@
Immediate from 3.2, 3.4, Lemma 3.6. Well-posedness of (17) is established in Theorem 3.9 below. ∎
@@ 3.3 Well posedness of the Volterra FBSDE
$\widehat{X}^{m,\xi}_{t}:=\mathcal{K}(0,t)\big({X}^{m,\xi}_{t}-M(m)_{t}\big),\hskip 2.84544pt\widehat{Y}^{m,\xi}_{t}:=\mathcal{K}(t,0)^{\top}{Y}^{m,\xi}_{t},$ (19)
$\widehat{\Sigma}_{t}:=\mathcal{K}(0,t)\Sigma_{t},\hskip 2.84544pt\widehat{Z}_{t}^{m,\xi}:=\mathcal{K}(s,0)^{\top}Z_{s}.$ (20)
@@
The first equation of the above system is a symmetric (see 3.8) Riccati matrix differential equation with terminal condition, and the second is a non-homogeneous linear matrix differential equation with terminal condition.
@@
which is an affine equation with respect to $\widehat{X}^{m,\xi}_{t}$ that obviously has a unique solution $\widehat{X}^{m,\xi}_{t}\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})$. Finally, one can directly verify via Itô’s formula (see for example [50, Corollary 6.4] ($V(t,x)\in C^{1,2}([0,T]\times\mathbb{R}^{d};\mathbb{R}^{d})$), and the terminal conditions for $P_{T},\Pi_{T}$, that $V(t,\widehat{X}^{m,\xi}_{t})$ satisfies the backward equation of (21). In other words, we have found a solution of (21) on the whole $[0,T]$, where the backward component is expressed via an affine decoupling field with the help of the forward component. Hence, to get uniqueness for the solution of (21) we apply [19, Proposition 4.8], and so we have arrived at the following.
@@
has a unique global solution on $[0,T]$. But this follows [^6] We remind again 3.8. from A.2 with the rough components equal to zero. ∎
@@ 4.1 Equivalence of Volterra fixed point step with Volterra FBODE
From 3.7, for every $m\in C([0,T];\mathbb{R}^{d})$ the LQ stochastic Volterra optimal control problem has a unique solution, and the optimal state satisfies, along with its adjoint state, the Volterra system
@@
First we assume that $F$ has a fixed point $m^{0}$. From the analysis before 4.1 we have
@@
For the opposite direction, let $(\bar{x}^{0},\bar{y}^{0})$ be a solution of (29). From 3.9 there exists a unique triplet $(X^{\bar{x}^{0}},Y^{\bar{x}^{0}},Z^{\bar{x}^{0}})\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{H}^{2}([0,T];\mathbb{R}^{d\times q})$ such that
@@
This last system has the form of (17) with $\xi=0,M(\cdot)=0,\Sigma_{t}=0,C_{t}=0,S_{t}=0,S=0.$ In view of 3.9 it has a unique solution which is equal to $(0,0,0)$. Thus, $\bar{x}^{0}$ is a fixed point of F and $(\bar{x}^{0},\bar{y}^{0})=\big(\bar{x}^{0},\mathbb{E}[{Y}^{\bar{x}^{0}}_{\cdot}]\big)$. ∎

### v2

@@ 3.1 Assumptions, state dynamics and cost functional
From this point on (and throughout the rest of the paper), we assume that $(\Omega,\mathcal{F},(\mathcal{F}_{t})_{t\in\mathbb{R}_{+}},\mathbb{P})$ is the usual augmentation of a filtration generated by a $q$-dimensional Brownian motion $W$ and a $\sigma$-algebra $\mathcal{F}_{0}^{0}$ independent of WW [^8] This assumption is made to avoid technicalities with martingale representation arguments that are used in the derivation of the Volterra FBSDEs.. From now on and for the rest of this section, we will work under the following assumptions:
@@
over all admissible controls $\alpha\in\widehat{\mathcal{A}}$, where
@@ 3.2 The stochastic maximum principle
$=$ $\penalty 10000\ R^{m}_{1}(X^{\alpha,m,\xi},(\delta X)^{h},\alpha,\varepsilon h)+R^{m}_{2}((\delta X)^{h},\varepsilon h)$
$=$ $\penalty 10000\ \varepsilon R^{m}_{1}(X^{\alpha,m,\xi},(\delta X)^{h},\alpha,h)+\varepsilon^{2}R^{m}_{2}((\delta X)^{h},h).$
@@
for some constant $C^{\prime\prime}$ depending on $\mathcal{K},A,B,T,Q_{t},\bar{Q}_{t},R_{t},Q,\bar{Q}.$ Hence,
@@
$\widehat{Y}^{\alpha,m,\xi}_{t}$ $\penalty 10000\ =\mathcal{K}(T,0)^{\top}\big((Q+\bar{Q})X_{T}^{\alpha,m,\xi}-\bar{Q}Sm_{T}\big)$
$\quad+\int_{t}^{T}\mathcal{K}(s,0)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{\alpha,m,\xi}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}\mathcal{K}(0,s)^{\top}\widehat{Y}^{\alpha,m,\xi}_{s}\Big)\mathrm{d}s$
$\quad-\int_{t}^{T}\widehat{Z}^{\alpha,m,\xi}_{s}\mathrm{d}W_{s}.$
The latter is a classical affine backward SDE with a unique solution
$(\widehat{Y}^{\alpha,m,\xi},\widehat{Z}^{\alpha,m,\xi})\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{H}^{2}([0,T];\mathbb{R}^{d\times q}),$
see for example [19, Theorem 4.23], hence, for every $\xi\in\mathbb{L}^{2}(\mathcal{F}_{0}),\alpha\in\widehat{\mathcal{A}}$ and $m\in C([0,T];\mathbb{R}^{d})$, (3.2) also has a unique solution
$({Y}^{\alpha,m,\xi},{Z}^{\alpha,m,\xi})\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{H}^{2}([0,T];\mathbb{R}^{d\times q}).$
@@
$(Y^{\alpha,m,\xi}_{T})^{\top}{(\delta X)}_{T}^{h}$ $\penalty 10000\ =\big(\mathcal{K}(T,0)^{\top}Y^{\alpha,m,\xi}_{T}\big)^{\top}\big(\mathcal{K}(0,T){(\delta X)}_{T}^{h}\big)$
$\penalty 10000\ =\int_{0}^{T}\big(\mathcal{K}(t,0)^{\top}Y^{\alpha,m,\xi}_{t}\big)^{\top}\mathrm{d}\big(\mathcal{K}(0,t){(\delta X)}_{t}^{h}\big)$ (14)
$\quad+\int_{0}^{T}\big(\mathcal{K}(0,t){(\delta X)}_{t}^{h}\big)^{\top}\mathrm{d}\big(\mathcal{K}(t,0)^{\top}Y^{\alpha,m,\xi}_{t}\big).$
@@
$\int_{0}^{T}\big(\mathcal{K}(0,t){(\delta X)}_{t}^{h}\big)^{\top}\mathrm{d}\big(\mathcal{K}(t,0)^{\top}Y^{\alpha,m,\xi}_{t}\big)$
$=$ $\penalty 10000\ \int_{0}^{T}\Big(\big({(\delta X)}_{t}^{h}\big)^{\top}A_{t}^{\top}Y^{\alpha,m,\xi}_{t}\Big)\mathrm{d}t$
@@
$\mathbb{E}\Big[(Y^{\alpha,m,\xi}_{T})^{\top}{(\delta X)}_{T}^{h}\Big]=\penalty 10000\ -\mathbb{E}\!\left[\int_{0}^{T}\big({(\delta X)}_{t}^{h}\big)^{\top}\big((Q_{t}+\bar{Q}_{t})X_{t}^{\alpha,m,\xi}-\bar{Q}_{t}S_{t}m_{t}\big)\mathrm{d}t\right]$
$\quad+\penalty 10000\ \mathbb{E}\!\left[\int_{0}^{T}\Big((Y^{\alpha,m,\xi}_{t})^{\top}B_{t}h_{t}\Big)\mathrm{d}t\right].$
@@
$\left\{\begin{aligned} {X}^{m,\xi}_{t}&\penalty 10000\ =\mathcal{K}(t,0)\xi+M(m)_{t}\\ &\quad\quad+\int_{0}^{t}\mathcal{K}(t,s)\big(A_{s}{X}^{m,\xi}_{s}-B_{s}R_{s}^{-1}B_{s}^{\top}Y^{m,\xi}_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\mathcal{K}(t,s)\Sigma_{s}\mathrm{d}W_{s},\\ Y^{m,\xi}_{t}&\penalty 10000\ =\mathcal{K}(T,t)^{\top}\big((Q+\bar{Q})X_{T}^{m,\xi}-\bar{Q}Sm_{T}\big)\\ &\quad\quad+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{m,\xi}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}Y^{m,\xi}_{s}\Big)\mathrm{d}s\\ &\quad\quad-\int_{t}^{T}\mathcal{K}(s,t)^{\top}Z^{m,\xi}_{s}\mathrm{d}W_{s}.\end{aligned}\right$. (18)
@@
$\left\{\begin{aligned} {X}^{\alpha,m,\xi}_{t}&\penalty 10000\ =\mathcal{K}(t,0)\xi+M(m)_{t}\\ &\quad\quad+\int_{0}^{t}\mathcal{K}(t,s)\big(A_{s}{X}^{\alpha,m,\xi}_{s}+B_{s}\alpha_{s}+C_{s}m_{s}\big)\,\mathrm{d}s+\int_{0}^{t}\mathcal{K}(t,s)\Sigma_{s}\mathrm{d}W_{s},\\ Y^{\alpha,m,\xi}_{t}&=\penalty 10000\ \mathcal{K}(T,t)^{\top}\big((Q+\bar{Q})X_{T}^{\alpha,m,\xi}-\bar{Q}Sm_{T}\big)\\ &\quad\quad+\int_{t}^{T}\mathcal{K}(s,t)^{\top}\Big((Q_{s}+\bar{Q}_{s})X_{s}^{\alpha,m,\xi}-\bar{Q}_{s}S_{s}m_{s}+A_{s}^{\top}Y^{\alpha,m,\xi}_{s}\Big)\mathrm{d}s\\ &\quad\quad-\int_{t}^{T}\mathcal{K}(s,t)^{\top}Z^{\alpha,m,\xi}_{s}\mathrm{d}W_{s}.\end{aligned}\right$.
@@
System (18), derived by (17) as explained in Remark 3.5, a priori describes a necessary condition for $J^{m,\xi}(\cdot)$ to have a local minimizer in $\widehat{\mathcal{A}}$. Below we show that in fact it provides a necessary and sufficient condition for $J^{m,\xi}(\cdot)$ to have a unique global minimizer in $\widehat{\mathcal{A}}$, as in the non-Volterra setting.
@@
Thus, modulo well-posedness of the Volterra FBSDE (18), we have showed the following:
@@
Immediate from Remarks 3.2, 3.4 and Lemma 3.6. Well-posedness of (18) is established in Theorem 3.10 below. ∎
Remark 3.8.
At this point let us note one of the main conceptual advantages from using the mild (or Volterra) formulation. Due to the nature of the cost function, $Y^{\alpha,m,\xi}_{t}$ was introduced as a solution of a tailor made Volterra BSDE. This allowed us to use the standard product formula for semimartingales to calculate the product $(Y^{\alpha,m,\xi}_{T})^{\top}{(\delta X)}_{T}^{h}$, see (14) above. If we worked directly in a rough setting the analog of $Y^{\alpha,m,\xi}_{t}$ would have been introduced as a solution of a rough BSDE, and at this step we would needed to use a rough product formula such as Theorem A.3. As seen therein, this requires strong integrability conditions for the diffusion that in general one cannot get from the martingale representation theorems. Here, in the linear-quadratic setting, we have available (global) decoupling fields, so we can guarantee these integrability conditions and offer an alternative route. But in more general settings [^10] Of course with affine dependence on state in the rough integral in order to be able to get the mild formulation of the dynamics., where (global) decoupling fields are not available, it is not clear if and how one can do this. However, even in these situations we could still push through with the mild formulation, the difference would be then that the stochastic system (18) above should be written in a hybrid form with forward rough and backward Volterra dynamics.
@@ 3.3 Well posedness of the Volterra FBSDE
$\begin{split}\widehat{X}^{m,\xi}_{t}&:=\mathcal{K}(0,t)\big({X}^{m,\xi}_{t}-M(m)_{t}\big),\\ \widehat{Y}^{m,\xi}_{t}&:=\mathcal{K}(t,0)^{\top}{Y}^{m,\xi}_{t},\\ \widehat{\Sigma}_{t}&:=\mathcal{K}(0,t)\Sigma_{t},\\ \widehat{Z}_{t}^{m,\xi}&:=\mathcal{K}(s,0)^{\top}Z_{s}.\end{split}$ (20)
@@
The first equation of the above system is a symmetric (see Remark 3.9) Riccati matrix differential equation with terminal condition, and the second is a non-homogeneous linear matrix differential equation with terminal condition.
@@
which is an affine equation with respect to $\widehat{X}^{m,\xi}_{t}$ that obviously has a unique solution $\widehat{X}^{m,\xi}_{t}\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})$. Finally, one can directly verify via Itô’s formula (see for example [51, Corollary 6.4] ($V(t,x)\in C^{1,2}([0,T]\times\mathbb{R}^{d};\mathbb{R}^{d})$), and the terminal conditions for $P_{T},\Pi_{T}$, that $V(t,\widehat{X}^{m,\xi}_{t})$ satisfies the backward equation of (21). In other words, we have found a solution of (21) on the whole $[0,T]$, where the backward component is expressed via an affine decoupling field with the help of the forward component. Hence, to get uniqueness for the solution of (21) we apply [19, Proposition 4.8], and so we have arrived at the following.
@@
has a unique global solution on $[0,T]$. But this follows [^11] We remind again Remark 3.9. from Theorem A.2 with the rough components equal to zero. ∎
@@ 4.1 Equivalence of Volterra fixed point step with Volterra FBODE
From Theorem 3.7, for every $m\in C([0,T];\mathbb{R}^{d})$ the LQ stochastic Volterra optimal control problem has a unique solution, and the optimal state satisfies, along with its adjoint state, the Volterra system
@@
First we assume that $F$ has a fixed point $m^{0}$. From the analysis before Remark 4.1 we have
@@
For the opposite direction, let $(\bar{x}^{0},\bar{y}^{0})$ be a solution of (29). From Theorem 3.10 there exists a unique triplet $(X^{\bar{x}^{0}},Y^{\bar{x}^{0}},Z^{\bar{x}^{0}})\in\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{S}^{2}([0,T];\mathbb{R}^{d})\times\mathbb{H}^{2}([0,T];\mathbb{R}^{d\times q})$ such that
@@
This last system has the form of (18) with $\xi=0,M(\cdot)=0,\Sigma_{t}=0,C_{t}=0,S_{t}=0,S=0.$ In view of Theorem 3.10 it has a unique solution which is equal to $(0,0,0)$. Thus, $\bar{x}^{0}$ is a fixed point of F and $(\bar{x}^{0},\bar{y}^{0})=\big(\bar{x}^{0},\mathbb{E}[{Y}^{\bar{x}^{0}}_{\cdot}]\big)$. ∎